Journal of Algebra, Vol. 111, No. 1, Nov 1987, 133–153.

Automorphism groups of Lorentzian lattices.

Richard E. Borcherds,

D.P.M.M.S., University of Cambridge, 16 Mill Lane, Cambridge, CB2 1SB, England. The study of automorphism groups of unimodular Lorentzian lattices I_{n,1} was started

by Vinberg. These lattices have an infinite reflection group (if n ≥ 2) and Vinberg showed that the quotient of the automorphism group by the reflection groups was finite if and only if n ≤ 19. Conway and Sloane rewrote Vinberg’s result in terms of the Leech lattice Λ, showing that this quotient (for n ≤ 19) was a subgroup of ·0 = Aut(Λ). In this paper we continue Conway and Sloane’s work and describe Aut(I_{n,1}) for n ≤ 23. In these cases there is a natural complex U associated to I_{n,1}, whose dimension is the virtual cohomological dimension of the “non-reflection part” G_{n} of Aut(I_{n,1}), and which is a point if and only if n ≤ 19. For n = 20, 21, and 22 the group G_{n} is an amalgamated product of 2 subgroups of ·0, while G_{23} is a direct limit of 6 subgroups of ·0. The group G_{24} seems to be much more complicated (although it would probably be just about possible to describe it). We also have a few results about G_{n} for large n; for example, if n is at least 18 and congruent to 2, 3, 4, 5, or 6 mod 8 then Aut(I_{n,1}) is a nontrivial amalgamated product. We find a few new lattices whose reflection group has finite index in the automorphism group; for example, the even sublattice of I_{21,1} of determinant 4.

1. Definitions.

In this section we summarize some standard definitions and results that we will use. Many of the results can be found in Serre [5] or Bourbaki [2].

We say that a group G acting on a set fixes a subset S if every element of G maps S into S.

A lattice L is a finitely generated free Z-module with an integer valued bilinear form, written (x, y) for x and y in L. The type of a lattice is even (or II) if the norm x^{2} = (x, x) of every element x of L is even, and odd (or I) otherwise. If L is odd then the vectors in L of even norm form an even sublattice of index 2 in L. L is called positive definite, Lorentzian, nonsingular, etc. if the real vector space L ⊗ R is.

If L is a lattice then L′ denotes its dual in L ⊗ R; i.e. the vectors of L ⊗ R which have integral inner products with all elements of L. The dual L′ contains L and if L is nonsingular then L′/L is a finite abelian group whose order is called the determinant of L. (If L is singular we say it has determinant 0.) The lattice L is called unimodular if its determinant is 1. If S is any subset of L then S^{⊥} is the sublattice of elements of L orthogonal to S.

Even unimodular lattices of a given signature and dimension exist if and only if there is a real vector space with that signature and dimension and the signature is divisible by 8. Any two indefinite unimodular lattices with the same type, dimension, and signature are isomorphic. I_{m,n} and II_{m,n} (m ≥ 1, n ≥ 1) are the unimodular lattices of dimension m + n, signature m − n, and type I or II.

A vector v in a lattice L is called primitive if v/n is not in L for any n > 1. A root of a lattice L is a primitive vector r of L such that reflection in the hyperplane r^{⊥} maps L to itself. This reflection maps v in L to v − 2r(v, r)/(r, r). Any vector r in L of norm 1 or 2 is a root.

If L is unimodular then there is a unique element c in L/2L such that (c, v) ≡ v^{2} mod 2 for all v in L. The vector c or any inverse image of c in L is called a characteristic vector of L, and its norm is congruent to the signature of L mod 8.

We now summarize some definitions and basic properties of finite root systems. “Root system” will mean “root system all of whose roots have norm 2” unless otherwise stated, so we only consider components of type a_{n}, d_{n}, e_{6}, e_{7}, e_{8}. We use small letters x_{n} to stand for spherical Dynkin diagrams. The types e_{3}, e_{4}, e_{5} are the same as a_{2}a_{1}, a_{4}, and d_{5}. The types d_{2} and d_{3} are the same as a_{1}^{2} and a_{3}.

The norm 2 vectors in a positive definite lattice A form a root system which we call the root system of A. The hyperplanes perpendicular to these roots divide A ⊗ R into regions called Weyl chambers. The reflections in the roots of A generate a group called the Weyl group of A, which acts simply transitively on the Weyl chambers of A. Fix one Weyl chamber D. The roots r_{i} which are perpendicular to the faces of D and which have inner product at most 0 with the elements of D are called the simple roots of D. (These have opposite sign to what are usually called the simple roots of D. This is caused by the irritating fact that the usual sign conventions for positive definite lattices are not compatible with those for Lorentzian lattices. With the convention we use something is in the Weyl chamber if and only if it has inner product at most 0 with all simple roots, and a root is simple if and only if it has inner product at most 0 with all simple roots not equal to itself.)

The Dynkin diagram of D is the set of simple roots of D. It is drawn as a graph with one vertex for each simple root of D and two vertices corresponding to the distinct roots r, s are joined by −(r, s) lines. (If A is positive definite then two vertices are always joined by 0 or 1 lines. We will later consider the case that A is Lorentzian and then its Dynkin diagram may contain multiple bonds, but these are not the same as the multiple bonds appearing in b_{n}, c_{n}, f_{4}, and g_{2}.) The Dynkin diagram of A is a union of components of type a_{n}, d_{n}, e_{6}, e_{7} and e_{8}. The Weyl vector ρ of D is the vector in the vector space spanned by roots of A which has inner product −1 with all simple roots of D. It is in the Weyl chamber D and is equal to half the sum of the positive roots of D, where a root is called positive if its inner product with any element of D is at least 0. A tip of a spherical Dynkin diagram is one of the points of weight 1. The number of tips of a_{n}, d_{n} (n ≥ 3), e_{6}, e_{7}, e_{8} is n, 3,2,1,0. The tips of a connected Dynkin diagram R are in natural 1:1 correspondence with the nonzero elements of the group A′/A, where A is the lattice generated by R.

The automorphism group of A is a split extension of its Weyl group by N, where N is the group of automorphisms of A fixing D. This group N acts on the Dynkin diagram of D and Aut(A) is determined by its Dynkin diagram R, the group N, and the action of N on R. There is a unique element i of the Weyl group taking D to −D, and −i is called the opposition involution of D and is denoted by σ or σ(D). The element σ fixes D and has order 1 or 2. (Usually −σ is called the opposition involution.)

If A is Lorentzian or positive semidefinite then we can still talk about its root system and A still has a fundamental domain D for its Weyl group and a set of simple roots. A may or may not have a Weyl vector.

We now describe the geometry of Lorentzian lattices and its relation to hyperbolic space.

Let L be an (n + 1)-dimensional Lorentzian lattice (so L has signature n − 1). Then the vectors of L of zero norm form a double cone and the vectors of negative norm fall into two components. The vectors of norm −1 in one of these components form a copy of n-dimensional hyperbolic space H_{n}. The group Aut(L) is a product Z_{2} × Aut_{+}(L), where Z_{2} is generated by −1 and Aut_{+}(L) is the subgroup of Aut(L) fixing each component of negative norm vectors. See Vinberg [8] for more details.

If r is any vector of L of positive norm then r^{⊥} gives a hyperplane of H_{n} and reflection in r^{⊥} is an isometry of H_{n}. If r has negative norm then r represents a point of H_{n} and if r is nonzero but has zero norm then it represents an infinite point of H_{n}.

The group G generated by reflections in roots of L acts as a discrete reflection group on H_{n} so we can find a fundamental domain D for G which is bounded by reflection hyperplanes. The group Aut_{+}(L) is a split extension of this reflection group by a group of automorphisms of D.

Finally we recall Conway’s calculation of Aut_{+}(II_{25,1}); see Conway [3] or Borcherds [1]. If Λ is the Leech lattice then II_{25,1} is isomorphic to the set of all points (λ, m, n) with λ in Λ, and m, n integers, with the norm given by (λ, m, n)^{2} = λ^{2} − 2mn. If w = (0, 0, 1) then the roots of II_{25,1} which have inner product −1 with w are the points (λ, 1, \frac{1}{2}λ^{2} − 1) which form a set of simple roots of a fundamental domain D of the reflection group of II_{25,1}, so that Aut_{+}(II_{25,1}) is a split extension of the reflection group by Aut(D), and Aut(D) is isomorphic to the group ·∞ of affine automorphisms of Λ, which is in turn a split extension of Z^{24} by Conway’s group ·0 = Aut(Λ).

2. Notation.

We define notation for the rest of this paper. L is II_{25,1}, with fundamental domain D and Weyl vector w. We identify the simple roots of D with the affine Leech lattice Λ.

R and S are two sublattices of II_{25,1} such that R^{⊥} = S, S^{⊥} = R, and R is positive definite and generated by a nonempty set of simple roots of D. The Dynkin diagram of R is the set of simple roots of D in R and is a union of a’s, d’s, and e’s. The finite group R′/R is naturally isomorphic to S′/S (in more than one way) because II_{25,1} is unimodular, and subgroups of R′/R correspond naturally to subgroups of S′ containing S (in just one way). We fix a subgroup G of R′/R and write T for the subgroup of S′ corresponding to it, so that an element s of S′ is in T if and only if it has integral inner product with all elements of G. It is T that we will be finding the automorphism group of in the rest of this paper. For each component R_{i} of the Dynkin diagram of R the nonzero elements of the group ⟨R_{i}⟩′/⟨R_{i}⟩ can be identified with the tips of R_{i}, and R′/R is a product of these groups. Any automorphism of D fixing R acts on the Dynkin diagram of R and on R′/R, and these actions are compatible with the map from tips of R to R′/R. In particular we can talk of the automorphisms of D fixing R and G. We will also write R for the Dynkin diagram of R. We write x′ for the projection of any vector x of II_{25,1} into S.

Example. Λ contains a unique orbit of d_{25}’s; let R be generated by a d_{n} (n ≥ 2) contained in one of these d_{25}’s. (N.B.: Λ contains two classes of d_{16}’s and d_{24}’s.) Then R^{⊥} = S is the even sublattice of I_{25−n,1} and R′/R has order 4. We can choose G in R′/R to have order 2 in such a way that T is I_{25−n,1}. We will use this to find Aut(I_{m,1}) for m ≤ 23.

We write Aut(T) for the group of automorphisms of T induced by automorphisms of II_{25,1}. This has finite index in the group of all automorphisms of T and is equal to this group in all the examples of T we give. This follows from the fact that if g is an automorphism of T fixing S and such that the automorphism of T/S it induces is induced by an automorphism of R (under the identification of G with S′/S) then g can be extended to an automorphism of II_{25,1}.

3. Some automorphisms of T.

The group of automorphisms of D fixing (R, G) obviously acts on T. In this section we construct enough other automorphisms of T to generate Aut(T) and in the next few sections we find how these automorphisms fit together. Recall that Λ is the Dynkin diagram of D and R is a spherical Dynkin diagram of Λ.

Let r be any point of Λ such that r ∪ R is a spherical Dynkin diagram and write R′ for R ∪ r. If g′ is any element of ·∞ such that σ(R′)g′ fixes r and R then we define an automorphism g = g(r, g′) of II_{25,1} by g = σ(R)σ(R′)g′. (Recall that σ(X) is the opposition involution of X, which acts on the Dynkin diagram X and acts as −1 on X^{⊥}.) These automorphisms will turn out to be a sort of generalized reflection in the sides of a domain of T.

Lemma 3.1.

(1) If g = g(r, g′) fixes the group G then g restricted to T is an automorphism of T. g fixes R and G if and only if σ(R′)g′ does.

(2) g fixes the space generated by r′ and w′ and acts on this space as reflection in r^{′⊥}.

(3) If g′ = 1 then g acts on T as reflection in r^{′⊥}. (g′ can only be 1 if σ(R′) fixes r.)

Proof. Both σ(R) and σ(R′)g′ exchange the two cones of norm 0 vectors in L and fix R, so g fixes both the cones of norm 0 vectors and R and hence fixes R^{⊥} = S. If g also fixes G then it fixes T as T is determined by G and S. σ(R) acts as −1 on G and fixes R, so g fixes R and G if and only if σ(R′)g′ does. This proves (1).

σ(R′)g′ fixes r and R and so fixes r′. σ(R) acts as −1 on anything perpendicular to R, and in particular on r′, so g(r′) = [σ(R)σ(R′)g′](r′) = −r′. If v is any vector of L fixed by g′ and v′ is its projection into T, then g(v) = σ(R)σ(R′)v so that g(v) − v is in the space generated by R′ and hence g(v′) − v′ is in the space generated by r′. As g(r′) = −r′, g(v′) is the reflection of v′ in r^{′⊥}. In particular if v = w or g′ = 1 then g′ fixes v so g acts on v′ as reflection in r^{′⊥}. This proves (2) and (3). Q.E.D.

Lemma 3.2.

(0) The subgroup of ·∞ = Aut(D) fixing (R, G) maps onto the subgroup of Aut_{+}(T) fixing w′.

(1) If σ(R′) fixes (R, G) then g(r, 1) acts on T as reflection in r^{′⊥} and this is an automorphism of T.

(2) If σ(R′) does not fix (R, G) then we define a map f from a subset of Aut(D) to Aut(T)

as follows:

If h in Aut(D) fixes all points of R′ then we put f(h) = h restricted to T.

If h′ in Aut(D) acts as σ(R′) on R′ then we put f(h′) = g(r, h′) restricted to T. (This is a sort of twisted reflection in r^{′⊥}.)

Then the elements on which we have defined f form a subgroup of Aut(D) and f is an isomorphism from this subgroup to its image in Aut(T). f(h) fixes r^{′⊥} and w′ while f(h′) fixes r^{′⊥} and acts on w′ as reflection in r^{′⊥}. (So f(h) fixes the two half-spaces of r^{′⊥} while f(h′) exchanges them.)

Proof. Parts (0) and (1) follow from 3.1.

If h in Aut(D) fixes all points of R′ then it certainly fixes all points of R and G and so acts on T. h also fixes w and w′.

If h′ acts as σ(R′) on R′ then σ(R′)g′ fixes all points of R, so by 3.1(1), g(r, h′) is an automorphism of T, and by 3.1(2), g(r, h′) maps w′ to the reflection of w′ in r^{′⊥}. (There may be no such automorphisms g′, in which case the lemma is trivial.) It is obvious that f is defined on a subgroup of Aut(D), so it remains to check that it is a homomorphism.

We write h, i for elements of Aut(D) fixing all points of R′ and h′, i′ for elements acting as σ(R′). All four of these elements fix R′ and so commute with σ(R′). h, i, σ(R′)h′, and σ(R′)i′ fix R and so commute with σ(R). σ(R)^{2} = σ(R′)^{2} = 1. Using these facts it follows that

f(hi) = hi = f(h)f(i)
f(hi′) = σ(R)σ(R′)hi′ = hσ(R)σ(R′)i′ = f(h)f(i′)
f(h′i) = σ(R)σ(R′)h′i = f(h′)f(i)
f(h′i′) = h′i′
= σ(R)^{2}σ(R′)^{2}h′i′
= σ(R)^{2}σ(R′)h′σ(R′)i′
= σ(R)σ(R′)h′σ(R)σ(R′)i′
= f(h′)f(i′)

so f is a homomorphism. Q.E.D.

4. Hyperplanes of T.

We now consider the set of hyperplanes of T of the form r^{′⊥}, where r is a root of II_{25,1} such that r′ has positive norm. These hyperplanes divide the hyperbolic space of T into chambers and each chamber is the intersection of T with some chamber of II_{25,1}. We write D′ for the intersection of D with T.

In the section we will show that D′ is often a sort of fundamental domain with finite volume. It is rather like the fundamental domain of a reflection group, except that it has a nontrivial group acting on it, and the automorphisms of T fixing sides of D′ are more complicated than reflections.

Lemma 4.1. D′ contains w′ in its interior and, in particular, is nonempty.

Proof. To show that w′ is in the interior of D′ we have to check that no hyperplane r^{⊥} of the boundary of D separates w and w′, unless r is in R. r is a simple root of D with −(r, w) = 1 so it is enough to prove that (r, ρ) ≥ 0, where ρ = w − w′ is the Weyl vector of the lattice generated by R. −ρ is a sum of simple roots of R, so (r, ρ) ≥ 0 whenever r is a simple root of D not in R because all such simple roots have inner product ≤ 0 with the roots of R. This proves that w′ is in the interior of D′. Q.E.D.

Lemma 4.2. The faces of D′ are the hyperplanes r^{′⊥}, where r runs through the simple roots of D such that r ∪ R is a spherical Dynkin diagram and r is not in R. In particular D′ has only a finite number of faces because R is not empty.

Proof. The faces of D′ are the hyperplanes r^{′⊥} for the simple roots r of D such that r^{′⊥} has positive norm, and these are just the simple roots of D with the property in 4.2. D′ has only a finite number of faces because the Leech lattice (identified with the Dynkin diagram of II_{25,1}) has only a finite number of points at distance at most \sqrt{6} from any given point in R. Q.E.D.

Lemma 4.3. If R does not have rank 24 (i.e., T does not have dimension 2) then D′ has finite volume.

Proof. D′ is a convex subset of hyperbolic space bounded by a finite number of hyperplanes, and this hyperbolic space is not one dimensional as T is not two dimensional, so D′ has finite volume if and only if it contains only a finite number of infinite points. R is nonempty so it contains a simple root r of D. The points of D′ at infinity correspond to some of the isotropic subspaces of II_{25,1} in r^{⊥} and D. The hyperplane r^{⊥} does not contain w as (r, w) = −1, so the fact that D′ contains only a finite number of infinite points follows from 4.4 below (with V = r^{⊥}). Q.E.D.

Lemma 4.4. If V is any subspace of II_{25,1} not containing w then V contains only a finite number of isotropic subspaces that lie in D.

Proof. Let II_{25,1} be the set of vectors (λ, m, n) with λ in Λ, m and n integers, with the norm given by (λ, m, n)^{2} = λ^{2} − 2mn. We let w be (0, 0, 1) so that the simple roots are (λ, 1, λ^{2}/2 − 1). As V does not contain w there is some vector r = (v, m, n) in V^{⊥} with (r, w) ≠ 0, i.e., m ≠ 0. We let each norm 0 vector z = (u, a, b) which is not a multiple of w correspond to the point u/a of Λ ⊗ Q. If z lies in V then (z, r) = 0, so (u/a − v/m)^{2} = (r/m)^{2}, so u/a lies on some sphere in Λ ⊗ Q. If z is in D then u/a has distance at least \sqrt{2} from all points of Λ (i.e., it is a “deep hole”), but as Λ has covering radius \sqrt{2} these points form a discrete set so there are only a finite number of them on any sphere. Hence there are only a finite number of isotropic subspaces lying in V and D. Q.E.D.

Remark. If w is in V then the isotropic subspaces of II_{25,1} in V and D correspond to deep holes of Λ lying on some affine subspace of Λ ⊗ Q. There is s universal constant n_{0} such that in this case V either contains at most n_{0} isotropic subspaces in D or contains an infinite number of them. If w is not in V then V can contain an arbitrarily large number of isotropic subspaces in D.

5. A complex.

We have constructed enough automorphisms to generate Aut(T), and the problem is to fit them together to give a presentation of Aut(T). We will do this by constructing a contractible complex acted on by Aut(T). For example, if this complex is one dimensional it is a tree, and groups acting on trees can often be written as amalgamated products.

Notation. We write Aut_{+}(T) for the group of automorphisms of T induced by Aut_{+}(II_{25,1}). D_{r} is a fundamental domain of the reflection subgroup of Aut_{+}(T) containing D′. The hyperbolic space of T is divided into chambers by the conjugates of all hyperplanes of the form r^{′⊥} for simple roots r of D.

Lemma 5.1. Suppose that for any spherical Dynkin diagram R′ containing R and one extra point of Λ there is an element of ·∞ acting as σ(R′) on R′. Then Aut_{+}(T) acts transitively on the chambers of T and Aut(D_{r}) acts transitively on the chambers of T in D_{r}.

Proof. By lemma 3.2 there is an element of Aut_{+}(T) fixing any face of D′ corresponding to a root r of D and mapping D′ to the other side of this face. Hence all chambers of T are conjugates of D′. Any automorphism of T mapping D′ to another chamber in D_{r} must fix D_{r}, so Aut(D_{r}) acts transitively on the chambers in D_{r}. Q.E.D.

D_{r} is decomposed into chambers of T by the hyperplanes r^{′⊥} for r a root of II_{25,1}. We will write U for the dual complex of this decomposition and U′ for the subdivision of U. U has a vertex for each chamber of D_{r}, a line for each pair of chambers with a face in common, and so on. U′ is a simplicial complex with the same dimension as U with an n-simplex for each increasing sequence of n + 1 cells of U. U is not necessarily a simplicial complex and need not have the same dimension as the hyperbolic space of T; in fact it will usually have dimension 0, 1, or 2. For example, if D_{r} = D′ then U and U′ are both just points.

Lemma 5.2. U and U′ are contractible.

Proof. U is contractible because it is the dual complex of the contractible space D_{r}. (D_{r} is even convex.) U′ is contractible because it is the subdivision of U. Q.E.D.

Theorem 5.3. Suppose that Aut(D_{r}) acts transitively on the maximal simplexes of U′, and let C be one such maximal simplex. Then Aut(D_{r}) is the sum of the subgroups of Aut(D_{r}) fixing the vertices of C amalgamated over their intersections.

Proof. By 5.2, U′ is connected and simply connected. C is connected and by assumption is a fundamental domain for Aut(D_{r}) acting on U′. By a theorem of Macbeth (Serre [6, p. 31]) the group Aut(D_{r}) is given by the following generators and relations:

Generators: An element ĝ for every g in Aut(D_{r}) such that C and g(C) have a point in common.

Relations: For every pair of elements (s, t) of Aut(D_{r}) such that C, s(C), and u(C) have a point in common (where u = st) there is a relation ŝt̂= û.

Any element of Aut(D_{r}) fixing C must fix C pointwise. This implies that C and g(C) have a point in common if and only if g fixes some vertex of C, i.e., g is in one of the groups C_{0}, C_{1}, … which are stabilizers of the vertices of C, so we have a generator ĝ for each g that lies in (at least) one of these groups. There is a point in all of C, s(C), and u(C) if and only if some point of C, and hence some vertex of C, is fixed by s and t. This means that we have a relation ŝt̂= û exactly when s and t both lie in some group C_{i}. This is the same as saying that Aut(D_{r}) is the sum of the groups C_{i} amalgamated over their intersections. Q.E.D.

Example. If C is one dimensional then Aut(D_{r}) is the free product of C_{0} and C_{1} amalgamated over their intersection. If the dimension of C is not 1 then Aut(D_{r}) cannot usually be written as an amalgamated product of two nontrivial groups.

6. Unimodular lattices.

In this section we apply the results of the previous section to find the automorphism group of I_{m,1} for m ≤ 23.

Lemma 6.1. Let X be the Dynkin diagram a_{n} (1 ≤ n ≤ 11), d_{n} (2 ≤ n ≤ 11), e_{n} (3 ≤ n ≤ 8), or a_{2}^{2} which is contained in Λ. Then any automorphism of X is induced by an element of ·∞. If X is an a_{n} (1 ≤ n ≤ 10), d_{n} (2 ≤ n ≤ 25, n ≠ 16 or 24), e_{n} (3 ≤ n ≤ 8), or a_{2}^{2} then ·∞ acts transitively on Dynkin diagrams of type X in Λ.

Proof. A long, unenlightening calculation. See section 9. Q.E.D.

Remark. ·∞ acts simply transitively on ordered a_{10}’s in Λ. There are two orbits of d_{16}’s and d_{24}’s (see section 9 and example 2 of section 8) and many orbits of a_{n}’s for n ≥ 11.

Notation. We take R to be a d_{n} contained in a d_{25} for some n with 2 ≤ n ≤ 23. If R is d_{4} we label the tips of R as x, y, z in some order, and if R is d_{n} for n ≠ 4 we label the two tips that can be exchanged by an automorphism of R as x and y. If n = 3 or n ≥ 5 we label the third tip of R as z. We let G be the subgroup of ⟨R′⟩/⟨R⟩ of order 2 which corresponds to the tip z if n ≥ 3 and to the sum of the elements x and y if n = 2. An automorphism of R fixes G if n ≠ 4 or if n = 4 and it fixes z. The lattice S = R^{⊥} is isomorphic to the even sublattice of I_{25−n,1}. We let T be the lattice corresponding to G that contains S, so that T is isomorphic to I_{25−n,1}.

Lemma 6.2. Any root r of V such that r ∪ R is a spherical Dynkin diagram is one of the following types: Type a: r ∪ R is d_{n}a_{1} (i.e., r is not joined to any point of R.) r′ is then a norm 2 vector

of T. Type d: r is joined to z if n ≥ 3 or to x and y if n = 2, so that r ∪ R is d_{n+1}. r′ has norm

1. Type e: r is joined to just one of x or y, so that r ∪ R is e_{n+1}. 2r′ is then a characteristic

vector of norm 8 − n in T.

Proof. Check all possible cases. Q.E.D.

In particular if r is of type a or d, or of type e with n = 6 or 7, then r′ (or 2r′) has norm 1 or 2 and so is a root of T. Note that in these cases σ(R ∪ r) fixes R and z and therefore G, so by 3.2(2), r^{′⊥} is a reflection of T. In the remaining four cases (r of type e with 2 ≤ n ≤ 5) σ(r ∪ R) does not fix both R and G.

Corollary 6.3. If n ≥ 6 then then reflection group of T = I_{25−n,1} has finite index in Aut(T). Its fundamental domain has finite volume and a face for each root r of Λ such that r ∪ R is a spherical Dynkin diagram.

Proof. This follows from the fact that all walls of D′ give reflections of T so D′ is a fundamental domain for the reflection group. By 4.3, D′ has finite volume. Q.E.D.

Remarks. The fact that a fundamental domain for the reflection group has finite volume was first proved in Vinberg [8] for n ≥ 8 and in Vinberg and Kaplinskaja [9] for n = 6 and 7 (i.e., for I_{18,1} and I_{19,1}). Conway and Sloane [4] show implicitly that for n ≥ 6 the non-reflection part of Aut(T) is the subgroup of ·∞ fixing R and their description of the fundamental domain of the reflection group in these cases is easily seen to be equivalent to that in 6.3. I_{18,1} and I_{19,1} have a “second batch” of simple roots of norm 1 or 2; from 6.2 we see that this second batch consists of the roots which are characteristic vectors and they exist because of the existence of e_{8} and e_{7} Dynkin diagrams. The non-reflection group of I_{20,1} is infinite because of the existence of e_{6} Dynkin diagrams and the fact that the opposition involution of e_{6} acts non-trivially on the e_{6}. (dim(I_{20,1}) = 1 + dim(II_{25,1}) + dim(e_{6}).)

From now on we assume that n is 2, 3, 4, or 5. Recall that D_{r} is the fundamental domain of the reflection group of T containing D and U is the complex which is the dual of the complex of conjugates of D in D_{r}.

Lemma 6.4. Aut(D_{r}) acts transitively on the vertices of U.

Proof. This follows from 6.1 and 5.1. Q.E.D.

Lemma 6.5. If n = 3, 4, or 5 then U is one dimensional and if n = 2, then U is two dimensional. (By the remarks after 6.3, U is zero dimensional for n ≥ 6.)

Proof. Let s and t be simple roots of Λ such that s ∪ R and t ∪ R are e_{n+1}’s, so that s^{′⊥} and t^{′⊥} give two faces of D′ inside D_{r}. Suppose that these faces intersect inside D_{r}. We have s^{′2} = t^{′2} = 2 − n/4, (s′, t′) ≤ 0, and r = s′ + t′ lies in T (as 2s′ and 2t′ are both characteristic vectors of T and so are congruent mod 2T). r^{′2} cannot be 1 or 2 as then the intersection of s^{′⊥} and t^{′⊥} would lie on the reflection hyperplane r^{⊥}, which is impossible as we assumed that s^{′⊥} and t^{′⊥} intersected somewhere in the interior of D_{r}. Hence

3 ≤ r^{2} = (s′ + t′)^{2} ≤ 2(2 − n/4) ≤ 2(2 − 2/4) = 3

so r^{2} = 3, n = 2, and (s′, t′) = 0.

If n = 3, 4, or 5 this shows that no two faces of D′ intersect in the interior of D_{r}, so the graph whose vertices are the conjugates of D′ in D_{r} such that two vertices are joined if and only if the conjugates of D′ they correspond to have a face in common is a tree. As it is the 1-skeleton of U, U must be one dimensional.

If n = 2 then it is possible for s^{′⊥} and t^{′⊥} to intersect inside D_{r}. In this case they must intersect at right angles, so U contains squares. However, in this case s and t cannot be joined to the same vertex of R = a_{1}^{2}, and in particular it is not possible for three faces of D′ to intersect inside D_{r}, so U is two dimensional. (Its two-dimensional cells are squares.) Q.E.D.

If X is a Dynkin diagram in Λ we will write G(X) for the subgroup of ·∞ fixing X. If X_{1} ⊂ X_{2} ⊂ X_{3} ⋯ is a sequences of Dynkin diagrams of Λ we write G(X_{1} ⊂ X_{2} ⊂ X_{3} ⋯) for the sum of the groups G(X_{i}) amalgamated over their intersections.

Theorem 6.6. The non-reflection part of Aut_{+}(T) = Aut_{+}(I_{25−n,1}) is given by

G(d_{5} ⊂ e_{6}) if n = 5,
G(d_{4}^{∗} ⊂ d_{5}) if n = 4 (where d_{4}^{∗} means that one of the tips of the d_{4} is labeled),
G(a_{3} ⊂ a_{4}) if n = 3,
G(a_{1}^{2} ⊂ a_{1}a_{2} ⊂ a_{2}^{2}) if n = 2.

Proof. By 6.4, Aut(D_{r}) acts transitively on the vertices of U. Using this and 6.1 it is easy to check that it acts transitively on the maximal flags of U, or equivalently on the maximal simplexes of the subdivision U′ of U. For example, if n = 5 this amounts to checking that the group G(d_{5}) acts transitively on the e_{6}’s containing a d_{5}.

By 5.3 the group Aut(D_{r}) is the sum of the groups fixing each vertex of a maximal simplex C of U′ amalgamated over their intersections. By 3.2 the subgroup of Aut(D_{r}) fixing the vertex D′ of U can be identified with G(R) = G(d_{n}) and the subgroup fixing a face of D′ can be identified with G(e_{n+1}) for the e_{n+1} corresponding to this face. These groups are the groups fixing two of the vertices of C, and if n = 2 it is easy to check that the group fixing the third vertex can be identified with G(a_{2}^{2}). Hence Aut(D_{r}) is G(d_{n} ⊂ e_{n+1}) if n = 3, 4, or 5 (where if n = 4 we have to use a subgroup of index 3 in G(d_{4})), and G(d_{2} ⊂ e_{3} ⊂ a_{2}^{2}) if n = 2. Q.E.D.

Examples. n = 3: Aut(I_{20,1}) The domain D′ has 30 faces of type a, 12 of type d, and 40 of type e. Aut(D′) is Aut(A_{6}) of order 1440 (where A_{6} is the alternating group of order 360). Aut(D_{r}) is

Aut(A_{6}) ∗_{H} (S_{3} ≀ Z_{2}),

where S_{3} ≀ Z_{2} = G(e_{6}) is a wreathed product and has order 72. H is its unique subgroup of order 36 containing an element of order 4. (Warning: Aut(A_{6}) contains two orbits of subgroups isomorphic to H. The image of H in Aut(A_{6}) is not contained in a subgroup of Aut(A_{6}) isomorphic to S_{3} ≀ Z_{2}.) Aut(D_{r}) has Euler characteristic 1/1440 + 1/72 − 1/36 = −19/1440.

n = 4: Aut(I_{21,1}). D′ has 42 faces of type a, 56 of type d, and 112 of type e. Aut(D′) is L_{3}(4) · 2^{2}. Aut(D_{r}) is

L_{3}(4).2^{2} ∗_{M_{10}} Aut(A_{6}).

Aut(A_{6}) has 3 subgroups of index 2, which are S_{6}, P SL_{2}(9), and M_{10}. Aut(D_{r}) has Euler characteristic −11/2^{8}.3^{2}.7.

n = 3: Aut(I_{22,1}). D′ has 100 faces of type a, 1100 of type d, and 704 of type e. Aut(D′) is HS.2, where HS is the Higman-Sims simple group, and Aut(D_{r}) is

HS.2 ∗_{H} H.2

where H is P SU_{3}(5) of order 2^{4}.3^{2}.5^{3}.7. Its Euler characteristic is 3.13/2^{10}.5^{5}.7.11.

n = 2: Aut(I_{23,1}). D′ has 4600 faces of type a, 953856 of type d, and 94208 of type e. Aut(D′) is ·2 × 2, where ·2 is one of Conway’s simple groups. Aut(D_{r}) is the direct limit of the groups

·2 × 2 = G(a_{1}^{2})
↗  ↖
McL ← P SU_{4}(3) → P SU_{4}(3).2
↓ ↓ ↓
G(a_{1}a_{2}) = McL.2 ← P SU_{4}(3).2 → P SU_{4}(3) · D_{8} = G(a_{2}^{2})

McL is the McLaughlin simple group and D_{8} is the dihedral group of order 8. The direct limit is generated by ·2 × 2 and an outer automorphism of McL; the P SU_{4}(3)’s are there to supply one additional relation. The Euler characteristic of Aut(D_{r}) is the sum of the reciprocals of the orders of the groups in the center and the vertices of the diagram above minus the sum of the reciprocals of the groups on the edges (see Serre [7]), which is 3191297/2^{20}.3^{6}.5^{3}.7.11.23.

Remark. For n = 2, 3, 4, 5, or 7 the number of faces of D′ of type e is (24/(n − 1) − 1)2^{12/(n−1)}, which is 23 · 2^{12}, 11 · 2^{6}, 7 · 2^{4}, 5 · 2^{3}, or 3 · 2^{2}. For n = 6 this expression is 20.06 … and there are 20 faces of type e. Table 3 of Conway and Sloane [4] gives the number of faces of type a and d of I_{m,1} for m ≤ 23.

7. More about I_{n,1}.

Here we give more information about I_{n,1} for 20 ≤ n ≤ 23. In the tables in Conway and Sloane [4] the heights of the simple roots they calculate appear to lie on certain arithmetic progressions; we prove that they always do. We then prove that the dimension of the complex U is the virtual cohomological dimension of the non-reflection part of Aut(I_{n,1}).

Notation. D is a d_{n} of Λ contained in a d_{25}. Let C be the sublattice of all elements of T = I_{25−n,1} which have even inner product with all elements of even norm. C contains 2T with index 2 and the elements of C not in 2T are the characteristic vectors of T. w′ is the projection of w into T and D_{r} is the fundamental domain of the reflection group of T containing w′.

Lemma 7.1. Suppose 2 ≤ n ≤ 5 Then all conjugates of w′ in D_{r} are congruent mod 2^{n−2}C. (If n ≥ 6 then D_{r} = D′ so there are no other conjugates of w′ in D_{r}.)

Proof. It is sufficient to prove that any two conjugates of w′ which are joined as vertices of the graph of D_{r} are congruent mod 2^{n−2}C because this graph is connected, and we can also assume that one of these vertices is w′ because Aut(D_{r}) acts transitively on its vertices. Let w′′ be a conjugate of w′ joined to w′.

w′′ is the reflection of w′ in some hyperplane e^{⊥}, where e is a characteristic vector of T of norm 8 − n, so e is in C. Therefore

w′′ = w′ − 2(w′, e)e/(e, e).

e = 2r′ for a vector r of Λ such that r ∪ R is an e_{n+1} diagram. The projections of r and w into the lattice I^{n} containing R are (\frac{1}{2}, \frac{1}{2}, …, \frac{1}{2}) and (0, 1, …, n − 1), which have inner product (n − 1)n/4. Hence (w′, r′) = (w, r) − (inner product of projections of w and r into ⟨R⟩) = −1 − (n − 1)n/4, so

−2(w′, e)/(e, e) = (4 + (n − 1)n)/(8 − n).

For 2 ≤ n ≤ 5 the expression on the right is equal to 2^{n−2}, so w′ and w′′ are congruent mod 2^{n−2}C. Q.E.D.

Theorem 7.2. Suppose n = 2, 3, 4, or 5 and let r be a simple root of the fundamental domain D_{r} of T = I_{25−n,1}.

If r^{2} = 1 then (r, w′) ≡ −n mod 2^{n−2}. If r^{2} = 2 then (r, w′) ≡ −1 mod 2^{n−1}.

Proof. (r, w′) = (s′, w′′) for some conjugate w′′ of w in D_{r} and some simple root s′ of I_{25−n,1} that is the projection of a simple root s of D into I_{25−n,1}. By 7.1, w′ is congruent to w′′ mod 2^{n−2}C so (r, w′) is congruent to (s′, w′) mod 2^{n−2} if r^{2} = 1 and mod 2^{n−1} if r^{2} = 2 (because elements of C have even inner product with all elements of norm 2). (s′, w′) is equal to (s, w)−(inner product of projections of s and w into R), which is −1 if s′ has norm 2 and −1 − (n − 1) if s′ has norm 1. Q.E.D.

This explains why the heights of the simple roots for I_{m,1} with 20 ≤ n ≤ 23 given in table 3 of Conway and Sloane [4] seem to lie on certain arithmetic progressions.

Now we show that the dimension of the complex U is the virtual cohomological dimension of Aut(D_{r}).

Lemma 7.3. The cohomological dimension of any torsion free subgroup of Aut(D_{r}) is at most dim(U).

Proof. Any such subgroup acts freely on the contractible complex U. Q.E.D.

Corollary 7.4. The virtual cohomological dimension of Aut(D_{r}) is equal to dim(U).

Proof. Aut(D_{r}) contains torsion-free subgroups of finite index, so by 7.3 the v.c.d. of Aut(D_{r}) is at most dim(U). For n ≤ 5, Aut(D_{r}) is infinite and so has v.c.d. at least 1, while for n = 2. Aut(D_{r}) contains subgroups isomorphic to Z^{2} (because there are 22dimensional unimodular lattices whose root systems generate a vector space of codimension 2) so Aut(D_{r}) has v.c.d. at least 2. Q.E.D.

Lemma 7.3 implies that Aut(D_{r}) contains no subgroups of the form Z^{i} with i > dim(U), and this implies that if L is a (24 − n)-dimensional unimodular lattice then the space generated by roots of L has codimension at most dim(U). This can of course also be proved by looking at the list of such lattices. (Vinberg used this in reverse: he showed that the non-reflection part of Aut(I_{m,1}) was infinite for m ≥ 20 from the existence of 19dimensional unimodular lattices with root systems of rank 18. There are two such lattices, with root systems a_{11}d_{7} and e_{6}^{3}; they are closely related to the two Niemeier lattices a_{11}d_{7}e_{6} and e_{6}^{4} containing an e_{6} component.)

8. Other examples.

We list some more examples of (not necessarily unimodular) Lorentzian lattices with their automorphism groups.

Example 1. R is an e_{8} in Λ so that T is II_{17,1}. The fundamental domain D′ of the reflection group has finite volume and its Dynkin diagram is the set of points of Λ not connected to e_{8}, which is a line of 17 points with 2 more points joined onto the 3rd and 15th points. This diagram was found in Vinberg [8].

Example 2. Similarly if R is one of the d_{16}’s of Λ not contained in a d_{17} then T is II_{9,1} and the points of Λ not joined to R form an e_{10} which is the Dynkin diagram of II_{9,1}.

Example 3. All e_{7}’s of Λ are conjugate; if R is one of them then T is the 19-dimensional even Lorentzian lattice of determinant 2. There are 3+21 roots r for which r ∪ R is a spherical Dynkin diagram and these 24 points are arranged as a ring of 18 points with an extra point joined on to every third point. The three roots joined to the e_{7} correspond to norm 2 roots r of T with r^{⊥} unimodular, while the other 21 roots correspond to norm 2 roots of T such that r^{⊥} is not unimodular. The non-reflection part of Aut_{+}(T) is S_{3} of order 6 acting in the obvious way on the Dynkin diagram.

Example 4. Let R be the unique orbit of e_{6}’s in Λ. Then T is the 20-dimensional even Lorentzian lattice of determinant 3. D′ is a fundamental domain for the reflection group of T and has 12 + 24 faces, coming from 12 roots of norm 6 and 24 of norm 2. The non-reflection group of Aut_{+}(T) is a wreath product S_{3} ≀Z_{2} of order 72. (Remark added in 1998: this example was first found by Vinberg in “The two most algebraic K3 surfaces”, Math. Ann. 265 (1983), no. 1, 1–21.)

Example 5. R is d_{4} and T is the even sublattice of I_{21,1} so that T has determinant 4. The domain D′ is a fundamental domain for the reflection group of T and has 168 walls corresponding to roots of norm 4 and 42 walls corresponding to roots of norm 2. Aut(D) is isomorphic to L_{3}(4).D_{12} of order 2^{8}.3^{3}.5.7. T is a 22-dimensional Lorentzian lattice whose reflection group has finite index in its automorphism group; I do not know of any other such lattices of dimension ≥ 21. (Remark added in 1998: Esselmann recently proved in “Über die maximale Dimension von Lorentz-Gittern mit coendlicher Spiegelungsgruppe”, Number Theory 61 (1996), no. 1, 103–144, that the lattice T is essentially the only example of a Lorentzian lattice of dimension at least 21 whose reflection group has finite index in its automorphism group.) T is contained in three lattices isomorphic to I_{21,1} each of whose automorphism groups has index 3 in Aut(T). However the reflection groups of these lattices do not have finite index in their automorphism groups.

Example 6. R is a_{1} and T is the 25-dimensional even Lorentzian lattice of determinant 2. This time D′ is not a fundamental domain for the reflection group. It has 196560 faces corresponding to norm 2 roots and 16773120 faces perpendicular to norm 6 vectors (which are not roots). However, the simplicial complex of T is a tree so Aut(D′) is (·0)∗_{(·3)}(2×·3), i.e., it is generated by ·0 and an element t of order 2 with the relations that t commutes with some ·3 of ·0. D′ has finite volume but if any of its 16969680 faces are removed the resulting polyhedron does not!

9. The automorphism groups of high-dimensional Lorentzian lattices.

Notation. L is II_{8n+1,1} (n ≥ 1) and D is a fundamental domain of the reflection group of L.

X is the Dynkin diagram of D. In this section we will show that if 8n ≥ 24 then Aut(L) acts transitively on many subsets of X, and use this to generalize some of the results of the previous sections to higher dimensional lattices.

Lemma 9.1. Let R be a spherical Dynkin diagram. Suppose that whenever R′ is a spherical Dynkin diagram in X which is isomorphic to R plus one point r there is an element g of Aut(D) such that gσ(R′) fixes R (resp. fixes all points of R).

For any map f : R ↦ X we construct (M, f′, C), where M is the lattice f(R)^{⊥}, f′ is the map from R′/R to M′/M such that f(r) ≡ −f′(r) mod L for r in R′/R, C is the cone of M contained in the cone of L containing D.

If f_{1}, f_{2} are two such maps then the images f_{1}(R), f_{2}(R) (resp. f_{1} and f_{2}) are conjugate under Aut(L, D) if the two pairs (M_{1}, f′_{1}, C_{1}), (M_{2}, f′_{2}, C_{2}) are isomorphic.

Proof. It is sufficient to show that a triple (M, f′, C) determines f(R) (resp. f) up to conjugacy under Aut(L, D). Given M and f′ we can recover L as the lattice generated by R ⊕ M and the elements r ⊕ f′(r) for r in R′. We have a canonical map from R to this L, so we have to show that the Weyl chamber D of L is determined up to conjugacy by elements of the group fixing R and M (resp. fixing M and fixing all points of R.) This Weyl chamber is determined by its intersection with R and M, and its intersection with R is just the canonical Weyl chamber of R. Its intersection with M is in the cone C and is in some Weyl chamber of the norm 2 roots of M. All such Weyl chambers of M in C are conjugate under automorphisms of L fixing M and all points of R, so we can assume that the intersection with M is contained in some fixed Weyl chamber W of M.

By 3.1 and the assumption on R all the Weyl chambers of L whose intersection with M is in W are conjugate under the group of automorphisms of L fixing R (resp. fixing all points of R) and hence (M, f′, C) determines f(R) (resp. f). Q.E.D.

Corollary 9.2. If R is e_{6}, e_{7}, e_{8}, d_{4}, or d_{m} (m ≥ 6) then two copies of R in X are conjugate under Aut(D) if and only if their orthogonal complements are isomorphic lattices.

Proof. If R′ is any Dynkin diagram containing R and one extra point then σ(R′) fixes R. The result now follows from 9.1. Q.E.D.

Remark. If L is II_{9,1} or II_{17,1} then Aut(D) is not transitive on d_{5}’s. σ(e_{6}) does not fix the d_{5}’s in e_{6}.

Lemma 9.3. Aut(D) is transitive on e_{8}’s. The simple roots of D perpendicular to an e_{8} form the Dynkin diagram of II_{8n−7,1} and the subgroup of Aut(D) fixing the e_{8} is isomorphic to the subgroup of Aut(II_{8n−7,1}) fixing a Weyl chamber.

Proof. The transitivity on e_{8}’s is in 9.2. The rest of 9.3 follows easily. Q.E.D.

Lemma 9.4. If 8n ≥ 24 then for any e_{6} in X there is an element of Aut(D) inducing σ(e_{6}) on it.

Proof. By 9.2, Aut(D) is transitive on e_{6}’s so it is sufficient to prove it for one e_{6}. It is true form 8n = 24 by calculation and using 9.3 it follows by induction for 8n > 24. Q.E.D.

Theorem 9.5. Classification of d_{m}’s in X.

(1) Aut(D) acts transitively on e_{6}’s e_{7}’s, and e_{8}’s in X. if 8n ≥ 24 then for any e_{6} there is an element of Aut(D) inducing the nontrivial automorphism of this e_{6}.

(2) For any m with 4 ≤ m ≤ 8n + 1 there is a unique orbit of d_{m}’s in X such that d_{m}^{⊥} is not unimodular, unless m = 5 and 8n = 8 or 16. Any automorphism of such a d_{m} is induced by an element of Aut(D) if and only if m ≤ 8n − 13.

(3) For any m with 16 ≤ 8m ≤ 8n there is a unique orbit of d_{8m}’s such that d_{8m}^{⊥} is unimodular, and these are the only d’s whose orthogonal complement is unimodular. There is no element of Aut(D) inducing the nontrivial automorphism of d_{8m}.

Proof. Part (1) follows from 9.2 and 9.4 because there is only one isomorphism class of lattices of the form e_{i}^{⊥} for i = 6, 7, 8. From 9.4, 9.2, and 9.1 it follows that two d_{m}’s of X are conjugate under Aut(D) if and only if their orthogonal complements are isomorphic, unless m = 5 and n ≤ 2. d_{m}^{⊥} is either the even sublattice of I_{8n+1−m,1}, or m is divisible by 8 and d_{m}^{⊥} is II_{8n+1−m,1}. In the second case we must have m ≥ 16 because if m was 8 the Dynkin diagram of (II_{8n+1−m,1})^{⊥} would be e_{8} and not d_{8}. This shows that there is one orbit of d_{m}’s unless 8|m, m ≥ 16 or m = 5, n ≤ 2 in which case there are two orbits.

If d_{m}^{⊥} is unimodular then d_{m} is contained in an even unimodular sublattice of L, and there are no automorphisms of this lattice acting non-trivially on d_{m}, so there are no elements of Aut(D) inducing a nontrivial automorphism of d_{m}.

If d_{m}^{⊥} is not unimodular then there is an element of Aut(D) inducing a nontrivial automorphism of d_{m} if and only if there is an automorphism of the Dynkin diagram of I_{8n+1−m,1} acting non-trivially on M′/M, where M is the sublattice of even elements of I_{8n+1−m,1}. There is no such automorphism of I_{k} for k ≤ 13 and there is such an automorphism for k = 14, so there is an element of Aut(D) inducing a nontrivial automorphism of d_{m} for m = 8n − 13 and there is no such element if m ≥ 8n − 12. If m < 8n − 13 then our d_{m} is contained in a d_{8n−13} so there is still a nontrivial automorphism of d_{m} induced by Aut(D). Finally, if m = 4 and 8n ≥ 24 then as in the proof of 9.4 we see that there is some d_{4} such that Aut(D) induces all automorphisms of d_{4}. As Aut(D) is transitive on d_{4}’s, this is true for any d_{4}. Q.E.D.

Lemma 9.6. If 2 ≤ m ≤ 11 and 8n ≥ 24 then for any a_{m} in L there is an element of Aut(D) inducing the nontrivial automorphism of a_{m}.

Proof. This is true for 8n = 24 by calculation. There is an element of Aut(D) acting as σ(a_{m}) on a_{m} if and only if there is an automorphism of the Weyl chamber of the lattice M = a_{m}^{⊥} which acts as −1 on M′/M. The lattice M is isomorphic to N ⊕ e_{8}^{n−3}, where N is a_{m}^{⊥} for some a_{m} in II_{25,1} and N has an automorphism of its Weyl chamber, so M has one too. Hence for m ≤ 11 there is an element of Aut(D) inducing σ(a_{m}) on a_{m}. Q.E.D.

Corollary 9.7. If 8n ≥ 24 and m ≤ 10 then Aut(D) is transitive on a_{m}’s in D.

Proof. It follows from 9.6 and 9.5 that if R′ is any spherical Dynkin diagram in X containing a_{m} and one extra point (so R′ is a_{m+1}, d_{m+1}, e_{m+1}, or a_{m}a_{1}) then there is an element g of Aut(D) such that gσ(R′) fixes a_{m}. Hence by 9.1, Aut(D) is transitive on a_{m}’s. Q.E.D.

Corollary 9.8. If n ≥ 20 and n ≡ 4, 5 or 6 mod 8 then the non-reflection part of Aut(I_{n,1}) can be written as a nontrivial amalgamated product.

Proof. The results of this section show that the analogue of 6.1 is true for II_{8i+1,1} for 8i ≥ 24. This is all that is needed to prove the analogue of 6.6. Q.E.D.

(If n ≥ 23 then the group cannot be written as an amalgamated product of finite groups.)

Remark. If n ≥ 10 and n ≡ 2 or 3 mod 8 and G is the subgroup of Aut(I_{n,1}) generated by the reflections of non-characteristic roots, then Aut(I_{n,1})/G is a nontrivial amalgamated product.

References.

1. R. E. Borcherds, The Leech lattice. Proc. Roy. Soc. London Ser. A 398 (1985), no. 1815, 365–376.

2. N. Bourbaki, “Groupes et algèbres de Lie”, Chaps. 4–6, Hermann, Paris, 1968.

3. J. H. Conway, The automorphism group of the 26-dimensional even unimodular Lorentzian lattice. J. Algebra 80 (1983), no. 1, 159–163.

4. J. H. Conway, N. J. A. Sloane, Leech roots and Vinberg groups. Proc. Roy. Soc. London Ser. A 384 (1982), no. 1787, 233–258.

5. J.-P. Serre, A course in arithmetic. Graduate Texts in Mathematics, No. 7. Springer-Verlag, New York-Heidelberg, 1973.

6. J.-P. Serre, “Trees”, Springer-Verlag, Berlin-New York, 1980. ISBN: 3-540-10103-9 7. J.-P. Serre, Cohomologie des groupes discrets. Prospects in mathematics (Proc. Sym-

pos., Princeton Univ., Princeton, N.J., 1970), pp. 77–169. Ann. of Math. Studies, No. 70, Princeton Univ. Press, Princeton, N.J., 1971.

8. È. B. Vinberg, Some arithmetical discrete groups in Lobačevskiĭ spaces. Discrete subgroups of Lie groups and applications to moduli (Internat. Colloq., Bombay, 1973), pp. 323–348. Oxford Univ. Press, Bombay, 1975.

9. Vinberg, Kaplinskaja, On the groups O_{18,1}(Z) and O_{19,1}(Z), Soviet Math. 19, No 1 (1978) 194-197.